Light travels from air (n = 1) into glass, where its velocity reduces to 2x108 m/s. What is the index of refraction for glass? Take speed of light in air to be 3x108 m/s
1
Expert's answer
2020-09-07T08:56:38-0400

As per the given question,

Refractive index of air "(\mu_1) =1"

Speed of light in air "(c)=3\times 10^8 m/sec"

Velocity of light in the medium "(V)=2\times 10^8 m/sec"

Refractive index of medium "(\mu_2)=?"

We know that "\dfrac{\mu_2}{\mu_1} =\dfrac{c}{v}"

Now, substituting the values in the above,

"\mu_2 =\dfrac{c}{v}\mu_1"

"\Rightarrow \mu_2 =\dfrac{3\times 10^8}{2\times 10^8}"

"\Rightarrow \mu_2 =1.5"